OpenStudy (anonymous):

condense the expression 1/5 log3 32- 2log3 x+ 1/2 log3 y

OpenStudy (anonymous):

\[1/5\log_{3} 32-2\log_{3} x+1/2\log_{3} y\]

OpenStudy (anonymous):

http://www.youtube.com/watch?v=o1AVbmS9AU8&feature=youtu.be this video may help you.... let me try to solve it for u tho

OpenStudy (anonymous):

\[\log_{3} (\frac{2\sqrt{y}}{x^2})\]

OpenStudy (anonymous):

\[\Large\log_{3} (\frac{2\sqrt{y}}{x^2})\]